Expand the binomial (1 - 2y)^5.
1
Expert's answer
2019-03-22T12:37:52-0400

Solution.

(1-2y)^5 = 1-10y+40y^2-80y^3+80y^4-32y^5

Answer:

(1-2y)^5 = 1-10y+40y^2-80y^3+80y^4-32y^5